1)For a monopolistic firm demand and total cost (TC) functions are as follows:
q = 230 – (1/2) P and TC = 20 + (1/2) q2
What is the equilibrium level of P (P*) that maximizes the profit of firm?

2)) If the total cost function (TC) of a competitive firm is TC = 20 + 0.2 Q2 and the price is (P) = 400 what is the equilibrium level of quantity (Q*) that maximizes profit?
1
Expert's answer
2020-07-27T09:45:41-0400

1)"TC=20+(1\/2)*(230-(1\/2)P)^2=20+(1\/2)*(52,900-230P+P^2\/4)=20+26,450-115P+P^2\/8=26,470-115P+P^2\/8"

"Profit=P*Q-TC"

"P(230-(1\/2)P)-26,470+115P-P^2\/8=0"

"345P-(5\/8)P^2-26,470=0"

"P^*(maxProfit)=-\frac{345}{-2*(5/8)}=276"

2) "q=30"

"TC=20+(1\/2)*30=35"

"Q^*=\frac{35-20}{0.2}=75"